OpenStudy (anonymous):

Finish writing the equation of the line y = 1/4x + ? , given the points (-12, 0) and (4, 4). Type only the number that is missing after the final + or - sign.

OpenStudy (anonymous):

This is solvable with the point-slope form of the equation for a line: y - y1 = m(x - x1) y - 4 = (1/4)(x - 4) y - 4 = (1/4)x - 1 y = (1/4)x + 3

OpenStudy (anonymous):

The last equation is in slope intercept and the number is : 3

OpenStudy (anonymous):

The slope was already given as 1/4, but also could have been calculated: (y2 - y1)/(x2 - x1) = m (4 - 0)/[4 - (-12)] = 4/16 = 1/4

OpenStudy (reemii):

in the equation \(y = (1/4) x + ....\), only one info is missing. Only one point is needed to find this value. As tcaroll010 said, it is 3. Replace y by 4 and x by 4 (we use (4,4)) and the equation must be ok. this will give you the value of "?" .

OpenStudy (anonymous):

btw, that "3" is the y-intercept and the line looks like:

OpenStudy (anonymous):

So 1/4?

OpenStudy (anonymous):

No, "3"

OpenStudy (anonymous):

"3" is the # you want. "1/4" is just additional incidental information.

OpenStudy (anonymous):

okay! so 1/4 x 3?

OpenStudy (anonymous):

No, just "3"

OpenStudy (anonymous):

3 goes where that ? is

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok my error

OpenStudy (anonymous):

my mistake - wow

OpenStudy (anonymous):

np. Your question was basically to replace the "?" in: y = 1/4x + ? So, you get: y = 1/4x + 3

OpenStudy (reemii):

\(y = (1/4)x + p\). you want to find \(p\). If (4,4) is on the line, then the equation is satisfied at this point. So the equation \(4 = (1/4)4 + p\) is correct. This shows that \(p=3\).